OpenStudy (anonymous):

Alanna can detail a car in 45 minutes. Robbie can detail a car in 30 minutes. If they work together, how many minutes does it take them to detail a car?

OpenStudy (anonymous):

want the quick snappo method?

OpenStudy (anonymous):

15 minutes 18 minutes 75 minutes 90 minutes

OpenStudy (anonymous):

Yes!!

OpenStudy (anonymous):

ok oh and i see you have answer choices, obvious the last two are rather silly because working together is evidently quicker than working alone

OpenStudy (anonymous):

snap method \[\frac{30\times 45}{30+45}\] compute that easy right?

OpenStudy (anonymous):

seems easy enough

OpenStudy (anonymous):

1350/ 75

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

it is notice i only used the numbers given a detailed explanation if you like, but that always works in this problem yes, 18

OpenStudy (anonymous):

No, I'm good:)

OpenStudy (anonymous):

lol fine

OpenStudy (anonymous):

Will you help with another?

OpenStudy (anonymous):

if i can sure

OpenStudy (anonymous):

wow i am sure i can come up with some answer but this question goes right in to the "bad math" pile maybe she scatters the leaves in one second!

OpenStudy (anonymous):

lollll :D

OpenStudy (anonymous):

i mean for real maybe she just comes along with a leave blower and blows them all to hell where does this question come from ?

OpenStudy (anonymous):

Summer math packet :)))

OpenStudy (anonymous):

heck i will make a guess that they want you to solve \[\frac{1}{15}+\frac{1}{x}=\frac{1}{20}\] which gives \(x=-60\) go with 60 and tell your teacher that satellite says they should be ashamed

OpenStudy (anonymous):

hahaha!!!! I'll tell him/her when I meet them next year hahah a

OpenStudy (anonymous):

in truth i have no idea what they want you to solve, so i would not bet more than $1.25 on my answer

OpenStudy (anonymous):

No biggie :) I trust you